two bodies are projected vertically upwards from one point with the same initial velocity vo.The second body is projected to second after the first.how long after will the bodies meet?

Grade:12th Pass

1 Answers

arun
8 years ago
It is obvious that the two bodies will meet when the first one will be returning towards ground while the second one still hasn't reached top most point of its flight.

Suppose above thing happens after 'T' seconds and at this instant both the bodies are at height 'h' above the point of projection. Apply second equation of motion at height 'h':

for first body, h= vT -1/2gT^2 .....(1)
for second body, h= v(T-t) -1/2gt^2 .....(2)

from (1) and (2),
vT - 1/2gT^2 = vT -vt -1/2gT^2 +gT*t -gt^2

=> T= 1/2t + v/g